# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ Acknowledgments
$\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq\Gamma_{1}^{n+1}\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}+\sum\limits_{i=0}^{n}\Gamma_{1}^{i}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]=\Gamma_{1}^{n+1}\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}+(I-\Gamma_{1})^{-1}(I-\Gamma_{1}^{n+1})\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (157)
@@
Also, substituting (171), (172) and assumption III.1 into the term $\Gamma_{1}^{n+1}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}$, we obtain
$\Gamma_{1}^{n+1}\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}\leq o(\frac{\mu}{B})$ (178)
@@
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]=\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq o(\frac{\mu}{B})+\left[\begin{array}[]{cc}-O(\frac{1}{\mu})&-O(1)\\ -O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{cc}-O(1)&-O(\mu)\\ -O(\mu^{2})&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]\leq\left[\begin{array}[]{c}O(\frac{\mu}{B})+O(\mu^{2})\\ O(\mu^{2})\end{array}\right]$ (189)
@@
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq\|\mathcal{V}\|^{2}\mathds{E}\|{\mathcal{V}^{\sf T}\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq O(\frac{\mu}{B})+O(\mu^{2})$ (190)
@@
We finally examine the bounds of the centralized method. As mentioned via (131), in the centralized method, $\check{\boldsymbol{w}}_{n}$ is always 00. Thus, we only need to analyze $\mathds{E}\|\bar{\boldsymbol{w}}_{n}\|^{2}$, for which according to (C), we have:
@@
By comparing (191) and (193), we see that the decentralized methods have extra $O(\mu^{2})$ noisy terms related to the network heterogeneity and graph structure.
Appendix D Proof for Lemmas III.6 and III.7: Upper bound for the fourth-order moment
@@
Then for $\check{\boldsymbol{w}}_{n-1}$, similar to the proof when analyzing the second-order error in Appendix C, we consider $t=\rho(P_{\alpha})$ which is the spectral radius of $P_{\alpha}$and obtain
@@
where $(a)$ follows from Jensen’s inequality, and we observe from (191) and (D) that
@@
As for the centralized method, we substitute (193) into (D) and iterate it, then we obtain
$\mathds{E}\|\bar{\boldsymbol{w}}_{n}\|^{2}{\leq}\left(\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}+O(\frac{\mu^{2}}{B})\right)^{n+1}\mathds{E}\|\bar{\boldsymbol{w}}_{-1}\|^{4}+\frac{1-\left(\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}\right)^{n+1}}{1-\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}}\times O(\frac{\mu^{3}}{B^{2}})$ (263)
@@
Appendix E Proof for Lemmas III.6 and III.7: Approximation error of the short-term model
The argument is similar to [6] except that we now focus on nonconvex (as opposed to convex) risk functions. To clarify how far the algorithm can escape from a local minimum $w^{\star}$, it is necessary to assess the size of the distance between $\boldsymbol{w}_{k,n}$ and $w^{\star}$ rather than upper bound it. However, the dependence of $\mathcal{H}_{n-1}$ on ${{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n-1}$ makes the analysis difficult. This motivates the short-term model in (48). Consider
@@
Similar to appendix C, we multiply both sides of (E) by $\mathcal{V}^{\sf T}$ so that it can be decomposed as
@@
which is positive-definite since it is the Hessian matrix of the global risk $J$ at a local minimum. Then we have
$\bar{\boldsymbol{z}}_{n}$ $=(I_{M}-\mu\bar{H})\bar{\boldsymbol{z}}_{n-1}-\mu\frac{1}{\sqrt{K}}\mathds{1}^{\sf T}\mathcal{H}\mathcal{V}_{\alpha}\check{\boldsymbol{z}}_{n-1}+\mu\frac{1}{\sqrt{K}}\mathds{1}^{\sf T}(\mathcal{H}_{n-1}-\mathcal{H})\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}$
$\check{\boldsymbol{z}}_{n}$ $=(\mathcal{P}_{\alpha}-\mu\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}\mathcal{H}\mathcal{V}_{\alpha})\check{\boldsymbol{z}}_{n-1}-\frac{1}{\sqrt{K}}\mu\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}\mathcal{H}\mathds{1}\bar{\boldsymbol{z}}_{n-1}+\mu\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}(\mathcal{H}_{n-1}-\mathcal{H})\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}$ (286)
@@
$=$ $(1-O(\mu))\mathds{E}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu)\mathds{E}\|\check{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu)\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}$ (287)
where (a) follows from Jensen’s inequality and (271).
@@
$=$ $(\rho(P_{\alpha})+O(\mu^{2}))\mathds{E}\|\check{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu^{2})\mathds{E}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu^{2})\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}$ (288)
where $(a)$ follows from Jensen’s inequality. By combining (E) and (E), we obtain the recursion associated with the size of $\mathds{E}\|\mathcal{V}_{\alpha}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}$:
$\mathds{E}\|\mathcal{V}_{\alpha}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n}\|^{2}\\ \|\check{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq\Gamma^{\prime}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}\\ \|\check{\boldsymbol{z}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\mu)\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\\ O(\mu^{2})\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\end{array}\right]$ (295)
@@
$\mathds{E}\|\mathcal{V}_{\alpha}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=(I-\Gamma^{\prime})^{-1}(I-\Gamma^{\prime n+1})\left[\begin{array}[]{c}O(\mu)\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\\ O(\mu^{2})\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\end{array}\right]$ (301)
@@
Then we substitute (D), (306), (311) and (312) into (301), and obtain
$\mathds{E}\left[\begin{array}[]{c}\|\check{\boldsymbol{z}}_{n}\|^{2}\\ \|\bar{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq$ $\left[\begin{array}[]{cc}O(\frac{1}{\mu})&O(1)\\ O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{cc}O(1)&-O(\mu)\\ -O(\mu^{2})&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\mu^{1+2\gamma})\\ O(\mu^{2+2\gamma})\end{array}\right]\leq O(\mu^{2\gamma})=O(\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4})$ (321)
@@
$\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}=\mathds{E}\|\mathcal{V}^{-\sf T}\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=O(\mu^{2\gamma})$ (322)
@@
Similarly, for the centralized method, we only need to analyze $\mathds{E}\|\bar{\boldsymbol{z}}_{n}\|^{2}$. Iterating (E) gives
@@
We substitute (193), (264), (312) into (328), and use (E) and (E), then we obtain
@@
Appendix F Proof for Theorem III.8
In this section, we derive closed-form expressions for the excess-risk performance of centralized, consensus and diffusion strategies over a finite time horizon $n\leq O(\frac{1}{\mu})$. Specifically, we verify how far the algorithms can escape from the local minimum $w^{\star}$ in terms of the risk function. To do this, we need to use the upper bounds in Lemmas III.6 and III.7, where centralized and decentralized methods have different expressions. For simplicity, we use $\gamma^{\prime}$ to unify the results of decentralized and centralized methods in Lemmas III.6 and III.7, namely,
@@
=$\overset{(b)}{=}\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{2}_{\bar{H}}\pm O(\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{3})$ (332)
@@
$\mathrm{ER}_{n}=\frac{1}{K}\sum\limits_{k}\mathrm{ER}_{k,n}\overset{(a)}{=}\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I\otimes\bar{H}}\pm O((\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4})^{\frac{3}{4}})\overset{(b)}{=}\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma^{\prime}})$ (333)
where $(a)$ follows from Taylor expansion and the following inequality:
@@
and $(b)$ follows from the following inequality which is similar to (E) and (326):
@@
$\leq\|\bar{H}\|^{2}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}+2\|\bar{H}\|\sqrt{\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}}$
@@
and also
@@
Note that for any two invertible matrices $X$ and $Y$, where $X=O(1)$ and $Y=o(1)$, i.e., $X$ dominates $Y$, we have
@@
Note that we have:
$\left|\mu\mathds{E}\left(\sum\limits_{i=0}^{n}\mathcal{C}^{i}\mathcal{A}_{2}d\right)^{\sf T}(I\otimes\bar{H})(\mathcal{C}^{n+1}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime})+\mathds{E}\|\mathcal{C}^{n+1}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|_{I\otimes\bar{H}}\right|$
$\leq\mu\mathds{E}\left\|\left(\sum\limits_{i=0}^{n}\mathcal{C}^{i}\mathcal{A}_{2}d\right)^{\sf T}(I\otimes\bar{H})\mathcal{C}^{n+1}\right\|\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|+\|\mathcal{C}^{n+1}\|^{2}\|\bar{H}\|\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|^{2}$
$\overset{(a)}{\leq}O(\mu)\times o(\sqrt{\frac{\mu}{B}})+o(\frac{\mu}{B})$
=$\overset{(b)}{=}o(\mu^{0.5(3+\eta)})+o(\mu^{1+\eta})$ (385)
where $(a)$ follows from Assumption III.1 and (381), and $(b)$ follows from (49), then we have:
@@
where $(h)$ follows from Assumption III.3, and $(g)$ follows from the following equality:

### v2

@@ Acknowledgments
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\Gamma_{1}^{n+1}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{-1}\|^{2}\\ \|\check{\boldsymbol{w}}_{-1}\|^{2}\end{array}\right]+\sum\limits_{i=0}^{n}\Gamma_{1}^{i}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]=\Gamma_{1}^{n+1}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{-1}\|^{2}\\ \|\check{\boldsymbol{w}}_{-1}\|^{2}\end{array}\right]+(I-\Gamma_{1})^{-1}(I-\Gamma_{1}^{n+1})\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (163)
@@
Also, according to assumption III.1, we have
$\mathds{E}\|\bar{\boldsymbol{w}}_{-1}\|^{2}=\mathds{E}\|\frac{1}{\sqrt{K}}\sum\limits_{k=1}^{K}\tilde{\boldsymbol{w}}_{k,-1}\|^{2}\leq\sum\limits_{k=1}^{K}\mathds{E}\|\tilde{\boldsymbol{w}}_{k,-1}\|^{2}\leq o(\frac{\mu}{B})$ (184)
$\mathds{E}\|\check{\boldsymbol{w}}_{-1}\|^{2}=\mathds{E}\|\mathcal{V}^{\sf T}_{\alpha}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}\leq\|\mathcal{V}^{\sf T}_{\alpha}\|^{2}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}\leq o(\frac{\mu}{B})$ (185)
Then with (177) and (178), we obtain
$\Gamma_{1}^{n+1}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{-1}\|^{2}\\ \|\check{\boldsymbol{w}}_{-1}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{c}o(\frac{\mu}{B})\\ o(\frac{\mu}{B})\end{array}\right]$ (190)
@@
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{c}o(\frac{\mu}{B})\\ o(\frac{\mu}{B})\end{array}\right]+\left[\begin{array}[]{cc}-O(\frac{1}{\mu})&-O(1)\\ -O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{cc}-O(1)&-O(\mu)\\ -O(\mu^{2})&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]\leq\left[\begin{array}[]{c}O(\frac{\mu}{B})+O(\mu^{2})\\ O(\mu^{2})\end{array}\right]$ (203)
@@
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq\|\mathcal{V}\|^{2}\mathds{E}\|{\mathcal{V}^{\sf T}\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\|\mathcal{V}\|^{2}\mathds{E}(\|\bar{\boldsymbol{w}}_{n}\|^{2}+\|\check{\boldsymbol{w}}_{n}\|^{2})\leq O(\frac{\mu}{B})+O(\mu^{2})$ (204)
@@
We finally examine the bounds of the centralized method. As mentioned via (123), in the centralized method, $\check{\boldsymbol{w}}_{n}$ is always 00. Thus, we only need to analyze $\mathds{E}\|\bar{\boldsymbol{w}}_{n}\|^{2}$, for which according to (B), we have:
@@
By comparing (205) and (207), we see that the decentralized methods have extra $O(\mu^{2})$ noisy terms related to the network heterogeneity and graph structure. According to (131), the extra $O(\mu^{2})$ noisy terms also relate to the consensus distance.
Appendix C Proof for Lemma III.5 and III.6: Upper bound for the fourth-order moment
@@
Then for $\check{\boldsymbol{w}}_{n-1}$, similar to the proof when analyzing the second-order error in Appendix B, we consider $t=\rho(P_{\alpha})$ which is the spectral radius of $P_{\alpha}$ and obtain
@@
where $(a)$ follows from Jensen’s inequality, and we observe from (205) and (C) that
@@
As for the centralized method, we substitute (207) into (C) and iterate it, then we obtain
$\mathds{E}\|\bar{\boldsymbol{w}}_{n}\|^{4}{\leq}\left(\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}+O(\frac{\mu^{2}}{B})\right)^{n+1}\mathds{E}\|\bar{\boldsymbol{w}}_{-1}\|^{4}+\frac{1-\left(\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}\right)^{n+1}}{1-\frac{(1+\mu L)^{4}}{(1-O(\mu))^{3}}}\times O(\frac{\mu^{3}}{B^{2}})$ (277)
@@
Appendix D Proof for Lemma III.5 and III.6: Approximation error of the short-term model
The argument is similar to [6] except that we now focus on nonconvex (as opposed to convex) risk functions. To clarify how far the algorithm can escape from a local minimum, it is necessary to assess the size of the distance between $\boldsymbol{w}_{k,n}$ and $w^{\star}$ rather than upper bound it. However, the dependence of $\mathcal{H}_{n-1}$ on ${{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n-1}$ makes the analysis difficult. This motivates the short-term model in (47). Consider
@@
Similar to appendix B, we multiply both sides of (D) by $\mathcal{V}^{\sf T}$ so that it can be decomposed as
@@
which is positive-definite since it is the Hessian matrix of the global risk $J$ at a local minimizer. Then we have
$\bar{\boldsymbol{z}}_{n}$ $=(I_{M}-\mu\bar{H})\bar{\boldsymbol{z}}_{n-1}-\frac{\mu}{\sqrt{K}}\mathds{1}^{\sf T}\mathcal{H}\mathcal{V}_{\alpha}\check{\boldsymbol{z}}_{n-1}+\frac{\mu}{\sqrt{K}}\mathds{1}^{\sf T}(\mathcal{H}_{n-1}-\mathcal{H})\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}$
$\check{\boldsymbol{z}}_{n}$ $=(\mathcal{P}_{\alpha}-\mu\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}\mathcal{H}\mathcal{V}_{\alpha})\check{\boldsymbol{z}}_{n-1}-\frac{\mu}{\sqrt{K}}\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}\mathcal{H}\mathds{1}\bar{\boldsymbol{z}}_{n-1}+\mu\mathcal{V}_{\alpha}^{\sf T}\mathcal{A}_{2}(\mathcal{H}_{n-1}-\mathcal{H})\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}$ (300)
@@
=$\overset{(b)}{=}$ $(1-O(\mu))\mathds{E}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu)\mathds{E}\|\check{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu^{1+2\gamma})$ (301)
where $(a)$ follows from Jensen’s inequality and (285), and $(b)$ follows from (C).
@@
$\leq$ $(\rho(P_{\alpha})+O(\mu^{2}))\mathds{E}\|\check{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu^{2})\mathds{E}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}+O(\mu^{2+2\gamma})$ (302)
where $(a)$ follows from Jensen’s inequality, and $(b)$ follows from (C). By combining (D) and (D), we obtain the recursion associated with the size of $\mathds{E}\|\mathcal{V}_{\alpha}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}$:
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n}\|^{2}\\ \|\check{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq\Gamma^{\prime}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n-1}\|^{2}\\ \|\check{\boldsymbol{z}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\mu^{1+2\gamma})\\ O(\mu^{2+2\gamma})\end{array}\right]$ (309)
@@
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n}\|^{2}\\ \|\check{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]=(I-\Gamma^{\prime})^{-1}(I-\Gamma^{\prime n+1})\left[\begin{array}[]{c}O(\mu^{1+2\gamma})\\ O(\mu^{2+2\gamma})\end{array}\right]$ (317)
@@
Then we substitute (322), (327) and (328) into (317), and obtain
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n}\|^{2}\\ \|\check{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq$ $\left[\begin{array}[]{cc}O(\frac{1}{\mu})&O(1)\\ O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{cc}O(1)&-O(\mu)\\ -O(\mu^{2})&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\mu^{1+2\gamma})\\ O(\mu^{2+2\gamma})\end{array}\right]\leq\left[\begin{array}[]{c}O(\mu^{2\gamma})\\ O(\mu^{2+2\gamma})\end{array}\right]$ (339)
@@
$\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}=\mathds{E}\|\mathcal{V}^{-\sf T}\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}\leq\|\mathcal{V}^{-\sf T}\|^{2}\mathds{E}(\|\check{\boldsymbol{z}}_{n}\|^{2}+\|\bar{\boldsymbol{z}}_{n}\|^{2})=O(\mu^{2\gamma})$ (340)
@@
Similarly, for the centralized method, we only need to analyze $\mathds{E}\|\bar{\boldsymbol{z}}_{n}\|^{2}$. Iterating (D) gives
@@
We substitute (207), (278), (328) into (346), and use (D) and (D), then we obtain
@@
Appendix E Proof for Theorem III.7
In this section, we derive closed-form expressions for the excess-risk performance of centralized, consensus and diffusion strategies over a finite time horizon $n\leq O(\frac{1}{\mu})$. Specifically, we verify how far the algorithms can escape from the local minimum $J(w^{\star})$ on average. To do this, we need to use the upper bounds in Lemmas III.5 and III.6, where centralized and decentralized methods have different expressions. For simplicity, we use $\gamma^{\prime}$ to unify the results of decentralized and centralized methods in Lemmas III.5 and III.6, namely,
@@
=$\overset{(b)}{=}{\frac{1}{2}}\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{2}_{\bar{H}}\pm O(\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{3})$ (350)
@@
$\mathrm{ER}_{n}=\frac{1}{K}\sum\limits_{k}\mathrm{ER}_{k,n}\overset{(a)}{=}\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I_{K}\otimes\bar{H}}\pm O((\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4})^{\frac{3}{4}})=\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I_{K}\otimes\bar{H}}\pm O(\mu^{1.5\gamma^{\prime}})$ (351)
where the $O(\mu^{1.5\gamma^{\prime}})$ term can be omitted according to (349), and $(a)$ follows from the Taylor expansion and the following inequality:
@@
Similar to (131), we first clarify the relationship between $\mathrm{ER}_{n}$ and the consensus distance. Recalling equality (105), we have
$\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I_{K}\otimes\bar{H}}=\left\|\left[\begin{array}[]{c}\bar{\boldsymbol{w}}_{n}\\ \check{\boldsymbol{w}}_{n}\end{array}\right]\right\|^{2}_{I_{K}\otimes\bar{H}}=\|\bar{\boldsymbol{w}}_{n}\|^{2}_{\bar{H}}+\|\check{\boldsymbol{w}}_{n}\|^{2}_{I_{K-1}\otimes\bar{H}}$ (355)
For the term of $\bar{\boldsymbol{w}}_{n}$, we have
$\|\bar{\boldsymbol{w}}_{n}\|^{2}_{\bar{H}}\overset{(a)}{=}\|\sqrt{K}\tilde{\boldsymbol{w}}_{c,n}\|^{2}_{\bar{H}}=K\|\tilde{\boldsymbol{w}}_{c,n}\|^{2}_{\bar{H}}=\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|_{I_{K}\otimes\bar{H}}$ (356)
where $(a)$ follows from (127), and the block vector ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}$ defined in (126) is a collection of $\tilde{\boldsymbol{w}}_{c,n}$. Also, for $\check{\boldsymbol{w}}_{n}$, we obtain
$\|\check{\boldsymbol{w}}_{n}\|^{2}_{I_{K-1}\otimes\bar{H}}\overset{(a)}{=}\|\check{\boldsymbol{w}}_{n}\|^{2}_{(V_{\alpha}^{\sf T}V_{\alpha})\otimes\bar{H}}=\|(V_{\alpha}\otimes I_{M})\check{\boldsymbol{w}}_{n}\|^{2}_{I_{K}\otimes\bar{H}}\overset{(b)}{=}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}_{I_{K}\otimes\bar{H}}$ (357)
where $(a)$ follows from (96), and $(b)$ follows from (129). Substituting (355), (356) and (357) into (351), we obtain
$\mathrm{ER}_{n}=\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I_{K}\otimes\bar{H}}\pm O(\mu^{1.5\gamma^{\prime}})=\frac{1}{2K}\mathds{E}\{\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|_{I_{K}\otimes\bar{H}}+\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}_{I_{K}\otimes\bar{H}}\}\pm O(\mu^{1.5\gamma^{\prime}})$ (358)
Thus, decentralized methods include the additional term caused by consensus distance, which further increases the escaping efficiency compared to the centralized method.
We next analyze the exact expression of $\mathrm{ER}_{n}$. To do this, we resort to the short-term model ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}$ which can well approximate the true model ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$. Basically, we have
$\mathrm{ER}_{n}=\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}_{I\otimes\bar{H}}\pm O((\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4})^{\frac{3}{4}})\overset{(a)}{=}\frac{1}{2K}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma^{\prime}})$ (359)
where $(a)$ follows from the following inequality which is similar to (D) and (344):
@@
$\leq\|\bar{H}\|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}+2\|\bar{H}\|\sqrt{\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}}$
@@
Note that for any $n$ and sufficiently small $\mu$, we have
$(I-\mu\bar{H})^{n+1}\leq I,\quad\;I-(I-\mu\bar{H})^{n+1}\leq I,\quad\;\mathcal{P}_{\alpha}^{n+1}\leq I,\quad\;I-\mathcal{P}_{\alpha}^{n+1}\leq I$ (406)
which means the right-hand matrices minus the left-hand side ones gives non-negative definite matrices. Thus we can write
$(I-\mu\bar{H})^{n+1}=O(1),\quad\;I-(I-\mu\bar{H})^{n+1}=O(1),\quad\;\mathcal{P}_{\alpha}^{n+1}=O(1),\quad\;I-\mathcal{P}_{\alpha}^{n+1}=O(1)$ (407)
with which and (402), we obtain
@@
where $O(1)$ means constant-bounded terms whose norms do not affected by small $\mu$.
Note also that for any two invertible matrices $X$ and $Y$, where $X=O(1)$ which is a constant and $Y=o(1)$ which is related to a small variable, i.e., $X$ dominates $Y$, we have
@@
where $(h)$ follows from Assumption III.3, and $(g)$ follows from (407) and the following equality: